Cramers rule Cant find this problem here
Solve the system of equation using Cramer's Rule if it is applicable. If Cramer's Rule is not applicable, say so.
Parenthesis 2x+3y=41
2x-3y=-13
A Parenthesis is only in the beginning then the two equations top and bottom..
here are the choices I need
A. x=7,y=9 B. x=-9,y=7 C. x=-7,y=-9 D. x=9,y=7
I think it is B is that correct?
Help

Found 2 solutions by KnightOwlTutor, nycsub_teacher:
Answer by KnightOwlTutor(293)

If you plug in the x and y values for b you will find it incorrect.
x=-9,y=7
2x+3y=41 -18+21=39
2x-3y=-13 -18-21=-39
2x+3y=41
2x-3y=-13
Add each column
4x+0=28
divide both sides by 4
x=7
Plug in the value of 7 for x to get the y value
14+3y=41
Subtract 14 from both sides
3y=27
divide both sides by 3
y=9
Plug in the values of x and y in the second equation
(2*7)-(3*9)=-13
-13=-13

Answer by nycsub_teacher(90)

Determinants can be used to solve a linear system of equations using Cramer’s Rule.
Cramer’s Rule is used for Two Equations in Two Variables, which is what you have: variables x and y.
2x+3y = 41
2x-3y = -13
To find the determinant, we use:
D = 2(-3) - 2(3)
D = -6 -6
D = -12
To find x use:
x = Dx/D
We already know D, right?
We need to find Dx.
We find Dx = 41(-3) - 3(-13)
Dx = -123 + 39
Dx = -84
We can now find x.
So, x = -84/-12
Then x = 7
Lastly, we need to find y.
y = Dy/D
We know D to be -12, right?
To find y, we need to first find Dy.
We can find Dy = 2(-13) - 2(41)
Dy = -26 - 82
Dy = -108
We can now find y.
y = -108/-12
y = 9
The solution to the above system of equations in two variables is the point
(7, 9)
The two equations given to you meet or cross each other at the point (7, 9) and so, this is why that particular point is the solution.
Understood?
The answer is choice (A).